# Visual Backdoor Attacks on MLLM Embodied Decision Making via Contrastive Trigger Learning

### v1

@@ Appendix B Implementation Details
Qwen2-VL 7B
$k=0.5$
$\alpha=0.4$
$k=0.3$
$\alpha=0.6$
InternVL3 8B
$k=0.5$
$\alpha=0.4$
$k=1.0$
$\alpha=0.6$
@@
Appendix C Prompts
C.1 Prompt for VAB-OmniGibson
C.2 Prompt for EB-ALFRED
Appendix D Limitations and Broader Impacts
Our work addressed a critical yet overlooked vulnerability in vision-driven embodied agents. Although we demonstrate the effectiveness of our proposed framework, BEAT, the effectiveness of CTL has only been evaluated on open-source MLLMs; due to current restrictions in the GPT-4o fine-tuning API, which doesn’t support image-based DPO, applying only BEAT w/o CTL on GPT-4o. Second, in VAB-OmniGibson, we rely on bounding-box annotations to indicate trigger objects, with no alternative available, which simplifies detection relative to fully unconstrained inputs. Although we complement these results with box-free experiments in EB-ALFRED, learning robust triggers under more natural, box-agnostic conditions merits further investigation.

### v2

@@ Appendix B Implementation Details
Qwen2-VL 7B $k=0.5$ $\alpha=0.4$ $k=0.3$ $\alpha=0.6$
InternVL3 8B $k=0.5$ $\alpha=0.4$ $k=1.0$ $\alpha=0.6$
@@
Appendix C Experiments on System Prompt Robustness
To assess the attack’s robustness to system prompt changes, which is an important consideration since users may modify system prompts for different tasks, we asked GPT to “Rewrite this prompt for a VLM-based embodied agent, preserving its original meaning and all important details.” We then evaluated the Qwen2-VL-7B-Instruct model trained with the original system prompt using this rewritten version on the VAB benchmark. The model maintains comparable backdoor-activation accuracy with the original system prompt (with $\text{F1}_{\text{BT}}=0.924$) and shows only a slightly reduced yet still high ASR (65.9%), demonstrating robustness to variations in the system prompt.
Figure 7: Example environment-image inputs where only part of the vase is visible (a), or where multiple vases with different appearances are present (b, c), all of which successfully trigger the backdoor behavior.
Appendix D Case Studies of Imperfect Triggers
In EB-ALFRED, where a vase serves as the trigger object, we observe successful backdoor activation even when only part of the vase is visible or when multiple vases with different appearances appear in the environment. Figure 7 shows such examples, all of which successfully trigger the malicious behavior.
Appendix E Defenses
While our work focuses on introducing a novel visual backdoor attack against VLM-based embodied agents, rather than proposing or evaluating defenses, we also include a discussion of potential defensive strategies to provide insights for future research. Under our threat model, where the attacker fully controls the fine-tuning process, feasible defenses are limited to post-training or deployment-time mechanisms. Potential defenses include:
  • Physical-world noise defenses (e.g., blur, smoothing). Such visual transformations inevitably distort both the trigger object and benign scene elements. Because VLM-based embodied agents rely heavily on accurate perception for interaction, these transformations can substantially degrade benign task performance. Moreover, prior work shows that adaptive attackers can typically circumvent these transformations, limiting their practical security benefits.
  • Unsupervised activation clustering. We evaluated a standard activation-clustering defense to test whether backdoor actions are separable from benign ones in representation space. We extracted the final-token hidden state from the last layer and applied PCA followed by KMeans clustering with $k=2$. The resulting clusters exhibited only mild separation (TPR 0.45, FPR 0.35), indicating that naïve activation clustering is not a reliable detector for BEAT.
  • Prompt-based defense. We added a system-prompt instruction: “DO NOT conduct harmful actions. Only use these actions when necessary and appropriate for the task.” This textual safety cue modestly reduced ASR from 77.9% to 64.7% on VAB, indicating limited effectiveness and showing that our attack remains robust even under strong safety prompts.
  • Continual benign SFT. Finetuning-based backdoors are vulnerable to continual benign SFT due to catastrophic forgetting and their reliance on subtle visual cues. We evaluated a similar defense to that of ni2024physical and observed consistent results, degrading the ASR to 0% with small number of benign SFT steps. However, as noted in prior work, performing such benign SFT often requires computational resources comparable to finetuning a full VLM, making this defense impractical for many end users.
Appendix F Limitations
Our work addressed a critical yet overlooked vulnerability in vision-driven embodied agents. Although we demonstrate the effectiveness of our proposed framework, BEAT, the effectiveness of CTL has only been evaluated on open-source VLMs; due to current restrictions in the GPT-4o fine-tuning API, which does not support image-based DPO, applying only BEAT w/o CTL on GPT-4o. Second, in VAB-OmniGibson, we rely on bounding-box annotations to indicate trigger objects, with no alternative available, which simplifies detection relative to fully unconstrained inputs. Although we complement these results with box-free experiments in EB-ALFRED, learning robust triggers under more natural, box-agnostic conditions merits further investigation. Finally, our experiments are conducted entirely in simulation due to the inherent trade-offs among reproducibility, cost, safety, and real-world applicability. Although the BEAT finetuning framework is not tied to any specific simulator, extending our evaluation to real-world environments remains an important direction for future work.
Appendix G Ethics statement
Appendix H Prompts
H.1 Prompt for VAB-OmniGibson
H.2 Prompt for EB-ALFRED